OpenStudy (anonymous):

Complete the equation -6x^2 + 15x -6 = -3(2x - 1)(x + __)

OpenStudy (hesan):

it is -2 Because as we multiply +3 by -2 this will give the required constant i.e -6

OpenStudy (anonymous):

Thanks, I appreciate the help.

OpenStudy (hesan):

You Welcome!!!

OpenStudy (anonymous):

let we call our anonymous :) (a) : \(-3(2x-1)(x+a)=(-6x+3)(x+a)=-6x^2-6ax+3x+3a\) \(=-6x^2+(-6a+3)x+3a\) We know that the coefficient of x is \(15\) so : \(-6a+3=15 <==>-6a=12 <==>a=-2\) Well, \(-6x^2 + 15x -6 = -3(2x - 1)(x +(-2))\)